Science Report – September 28th

Crew 228 Science Report 28Sep2021
Crew Scientist / GreenHab Officer Inga Popovaite
Group Processes Study:
Continuing data collection for the study (University of Iowa IRB# 201911141) on small group interaction in space analog environments.
Other research projects:
Nothing to report.

Inga Popovaitė,


Science Report Apr 14th

Science Report 14-Apr-2021

· Extremophiles Research

o Collected 2 samples while on EVA, and samples stored in fridge in the Science Dome. Intended to collect 4 samples, so at a later date, a following EVA might need to return to the coordinates of this collection site.

· Hydroponic Garden

o Significant leaking from the compost and H2O (control) pipes. Monitoring with catch buckets underneath. Still unable to measure evaporation due to leak.

o Began to collect data on stalk height and root growth

o Green Hab heater to be turned on at approx. 5pm every evening with temperature set to 65F.

In addition, after this report is sent, Crew Botanist will be planting seeds this evening.

Olivia Ettlin & Julio Hernandez

Science Report – April 13th

Crew Scientist:

· Hydroponic garden completely assembled.

o Aeration system and pipes assembled and transported from RAM to the GreenHab

o Materials used from the RAM: gorilla glue, dewalt drill, hammer, razor blade, and needle-nosed pliers.

o Thermostat in the GreenHab set to 65 degrees F near 8:30pm

· Experimental steps completed today

o Calculated the amt of additives to be input into system.

o Administered additives (experimental variables: worm castings, BioLive fertilizers, compost biproducts)

· Troubleshooting

o Small leak at the ends of the pipes. Need to seal with caulking/ PVC pipe glue on Sol 4. Rate of evaporation may no longer be a feasible component to track unless leak can be sealed.

Crew Botanist:
Started to set up a botany experiment investigating the effect of different soil compositions on the growth of plants. The crew botanist recognized that there was a miscalculation for determining the group and population sizes. Namely, the weight fraction was used to calculate the quantities of different soil compositions instead of the volume fraction. If the problem was not corrected, the experimental group with increasing amount of martian regolith will decrease the overall volume of the "soil/regolith"; this would have meant that there would have been less material surrounding the seeds (and less potential nutrients). The crew botanist started to implement the fix but the completion of the work is rescheduled for tomorrow.

Red Planet People – MDRS Crew 245 "Team Patamars"To Mars and Beyond – For All!

Science Report – March 10th

Research report 10 March 2020

Crew 223

Submitted by Crew Scientist Marion Lebrun

1. LOAC and Weather Station:

Nothing to report. Everything is working as it should.

2. MegaARES:

Nothing to report. Everything is working as it should. We are waiting a bit longer to collect another round of data.

3. Solar oven:

The oven was taken outside a second time on sol 8. As the sun was not very bright, it only heated the 600mL of water from 12°C to 45°C during the EVA time of the day. Although the water was not as warm as the first time (due to the lack of sun light), Florian is satisfied with the collected data and the photos he took of the experiment.

4. BackPack and LCVG:

Nothing to report.

5. Foreign languages (Human Factors):

The experiment is now over (it ended on sol 7). I have noticed remarkable progress in communication within the pairs. All three teams managed to do their best performances on the last day although the LEGO model to build was the most complicated of all seven. When we come back to France, the videos and pictures taken will be sent to the Human Factors department of our school to be thoroughly analyzed.

6. Group collaboration (Human Factors):

The experiment started on sol 8. Two groups of 3 crew members play “Keep Talking and Nobody Explodes”: they have 5 minutes to defuse a bomb placed on a Martian rover. Two experts have the bomb defusing manual and need to communicate the information to the defuser who only sees the bomb screen. Just like the first experiment, this one aims at studying the communication within the teams, between the defuser and the experts. It is too early to collect relevant results.

7. Crew members’ individual planning display:

The display of the sol schedules and the hyperlink to access the available Internet data are now ready but we still cannot find a way to display a hyperlink giving direct access to Mission Support emails. This is the last thing Blandine and I would like to achieve by the end of the mission.

8. Relaxation and positive psychology (Human Factor):

Everything works well. The whole crew has been getting used to our daily relaxation time, it truly regulates our mood and gives us a bowl of positive energy to finish the day.

9. Music for plants (GreenHab):

Nothing to report. It is still too early to notice remarkable results.

10. Spirulina growth (GreenHab):

Nothing to report. Everything is working as it should, nothing new.

11. Aerospring (GreenHab):

Nothing to report.

12. Water consumption monitoring:

Aurelien realized what was wrong with the monitoring: it is the way the buttons that classify the different water usages are coded that needs to be changed. Other than that, everything else is working.

13. Supernova research:

No new observation due to bad weather but Florian still hopes the weather will improve in the next few days.

Science Report – March 7th

Research report 6 March 2020
Crew 223
Submitted by Crew Scientist Marion Lebrun

1. LOAC and Weather Station:
Nothing to report. Everything is working as it should. We hope the
instruments resist the heavy wind of today.

2. MegaARES:
Nothing to report. Everything is working as it should. We are waiting a
bit longer to collect another round of data.

3. Solar oven:
The oven was taken outside for the first time this morning during the
EVA. It managed to heat 600mL of water from 10°C to 67°C in a bit less
than two hours. Florian is very satisfied with this first round of data.
He now plans to repeat the experiment with the same amount of water but
with a different orientation towards the Sun to compare the yield of the
solar oven.

4. BackPack and LCVG:
Nothing to report.

5. Foreign languages (Human Factors):
Communication within the teams is well established and much more
efficient than on the first session. There are almost no mistakes on the
parts chosen to build the given models which shows how communication has
improved. I have noticed a lot less silents during the 10min of LEGO
building. Crew members ask for a lot more precision and details:
sometimes it might slow down the construction but it certainly avoids
misinterpretations. I am overall very satisfied with the way the
experiment is going.

6. Group collaboration (Human Factors):
Nothing to report.

7. Crew members’ individual planning display:
A first draft has been successfully tested although emails and Internet
data are still missing (but the current sol schedule with the sport
schedule is there). Blandine and I have decided to give up on the email
window being refreshed every hour; we are thus going to create a
hyperlink to the Mission Support email page that would be a shorter path
than having a crew member open his/her computer to check it. The same
thing will be done for the available Internet data.

8. Relaxation and positive psychology (Human Factor):
Nothing to report. Everything works well.

9. Music for plants (GreenHab):
Yesterday was the first session of the experiment: the radish plants
are stimulated by UV light and music (sequence of notes that favors the
synthesis of a protein making the plant UV resistant) for 7 minutes
every day.

10. Spirulina growth (GreenHab):
Nothing to report. Everything is working as it should, nothing new. It
is too early analyse the impact of the type of fertilizer on the
spirulina growth.

11. Aerospring (GreenHab):
We have increased the watering (20 to 25 minutes twice a day instead of
15 minutes twice a day) due to the very dry atmosphere. Hopefully it
will increase the lettuce growth.

12. Water consumption monitoring:
Nothing to report.

13. Soundless EVA:
We want to test a protocol for a soundless EVA. Our aim is to avoid any
verbal communication between the Habcom and the EVA team, and also
within the EVA team. Of course each of us will have a working radio in
case of a problem or emergency. Here is the detailed protocol :
– we stop communications as soon as we enter in the airlock.
– we use a white board and a marker to write down all the requested
numbers : rovers’ hours, levels of charge, static tank level and time.
If needed, this board might also be used to communicate between us.
– when the airlock light is switched off by the Habcom, it means that
the depressurization is finished and that we are allowed to go outside.
– once outside, each of us will have a specific role. Three will do the
usual checks while one (the communicator) will hold the board and meet
the others to allow them to write down their data. Then the communicator
will communicate those data (using signs and gestures we have chosen) to
the Habcom, who will be looking at him through the upper-deck window.
– then the EVA team will wait for a signal from the Habcom (still
through the window) giving them the authorization to leave the Hab zone.
– in each buddy team, one will guide the pilot for the U turn.
– before the EVA, we will decide which rover will be first to leave.
– during the rover ride, the first rover needs to know if the 2nd rover
is following. Thus we have imagine a specific protocol. Every 1 min
(signal will be given by the copilot who will have a watch), the first
rover will stop on the right side of the road. The 2nd rover will slowly
pass the first one by the left. Then the stopped rover will move and
follow the new first one. And so one. Of course we will drive slowly,
especially during passing.
– when we arrive at our target destination, the first rover stops in the
middle of the road for about 20sec (to allow the 2nd one to stop
behind). Then the 1st rover will move and park on the left side of the
road (not right side to avoid any misunderstanding with the stop on the
right, that is dedicated to let the 2nd rover pass). The communicator
will come to the 2 pilots with the board to let them right down the
level of charge of the rovers.
– when we leave our target destination, after the U turn, the 1st rover
stops on the right of the road and the 2nd rover passes the 1st one by
the left.
– during the exploration, we will use most of the gestures we already
use to communicate during our EVAs. We will of course stick together,
the first walker will often turn around to check that everybody is still
there and fine. Every 10min, the leader will ask, using a gesture, how
the team feels (regarding their suits and their fatigue). Each member
(including the leader) must answer on a scale from 1 to 5. If 1,
everything is fine; if 4, we go back home; if 5, we use the radio and go
back home. Of course, once again, each team member is free to use its
radio if he/she feels he/she needs to.
– when we arrive in the Hab zone, the challenge is to get the approval
from the Habcom that allows us to park near the Hab. The EVA team will
stop near the Science Dome to wait for the approval. During the last
30min of the EVA, the Habcom will check every 5min or so if heshe sees
the EVA team waiting for the approval. He/She will allow us to park with
a sign through the window.
– the communicator will once again communicate rover’s data to the
Habcom. Once received the Habcom will allow with a gesture the EVA team
to enter in the airlock.
– the last gesture will be from the EVA team, to inform the Habcom
through the small window of the airlock that the airlock is closed. Then
the Habcom will start the pressurisation by switching on the airlock’s
In our opinions, this protocol has a real interest in case of the loss
of the 4 radios of the EVA team.
Here is Luc’s feedback after having applied this protocol :
He made an extensive briefing before entering in the airlock to make
sure everybody (including the Habcom of course) had the protocol in mind.
The five minutes of depressurization were the opportunity for each of us
to think about their upcoming actions and tasks. We walked out of the
airlock, Aurélien installed the solar oven, while Clément and Marion
checked the 3 rovers, and Luc checked the level of water in the static
tank. Then Marion, as communicator, was in charge to communicate through
the window all our data to Valentin, the Habcom. She did it
successfully, as Valentin understood all the figures.
The rover ride was the opportunity to test the passing procedure. For
the EVA Leader Luc, it was the most challenging part of the protocol as
the pilot of the first rover has no visibility. But everything went
smooth and we can assert that this passing procedure is very efficient
and safe for both rovers.
The exploration was the chance to test our communication protocol,
especially the 1 to 5 scale, used to know how each EVA team member
feels. It was efficient, but we found that it can be improved. Because
two members answered 3 out of 5 for the fatigue but during the
debriefing, we noticed that they had a really different interpretation
of that number (one was about to ask to go back home, while the other
one was almost ok). So we need to clarify what each number corresponds
to, to make sure we all have the same reference scale.
For our return, the Habcom was waiting for us, looking through the
window with binoculars. He gave us the approval to park the rovers near
the RAM. Marion once again was in charge of communicating the rovers’
data to the Habcom. Then we entered in the airlock at 11:15 AM for the 5
last minutes of silence of the day !
It was an interesting experience for all of us, and it was the
opportunity to demonstrate that this type of protocol is efficient and
usable in case of loss of all our radios during an EVA !

14. Supernova research:
It is a project common with the Crew 222. Based on a file gathering
galaxies potentially having supernova, the two astronomers have selected
around 50 galaxies and the have been doing observations every week. For
now, they have not found anything but Florian keeps looking.

Science Report – March 4th

Research report 4 March 2020
Crew 223
Submitted by Crew Scientist Marion Lebrun

1. LOAC and Weather Station:
Everything is working as it should. SOL 0 was very windy and both experiments remained standing. Therefore, with the strong gusts of wind, the collected data should be interesting to analyse.

2. MegaARES:
Everything is working as it should. We have collected a first round of data; we have now 2.4Gb of Earth electric field data which is very satisfying.

3. Solar oven:
The oven is built (still in the Hab) and last night Florian downloaded the programs needed for the sensors. The sensors need to be tested before setting the oven on Mars.

4. BackPack and LCVG:
Nothing to report.

5. Foreign languages (Human Factors):
3 pairs have been formed : Aurélien and Florian speaking Spanish, Blandine and Clément speaking English, Valentin and Luc speaking German. Although speaking in a foreign language might have been disturbing at first for some of the crew members, after 2 days of experiment (SOL 1 and SOL 2), progress is definitely noticeable. Communication is clearly improving between the two members of each team; it was more efficient on SOL 2 day and the teams managed to build assemble more parts of LEGO than they did on SOL 1 – which is very encouraging.

6. Group collaboration (Human Factors):
Nothing to report.

7. Crew members’ individual planning display:
Blandine and Marion have decided what needs to be displayed: a dynamic schedule of the current sol, a window with Mission Support email refreshed every hour, the sport schedule for every member of the crew, and the available Internet data remaining. Blandine and Marion are
finishing their training on html before starting to actually code the app.

8. Relaxation and positive psychology (Human Factor):
Relaxation seems to have a greater impact on our crew than it had on Crew 222. As for positive psychology, the technique seems to be strengthening our crew’s cohesion.

9. Music for plants (GreenHab):
It is very hard to grow seeds in vitro since the atmosphere is too dry. This is why Crew 222 GreenHab Officer failed to grow her seeds in gelatin and why Valentin also failed to grow his on humidified paper towel. Valentin has thus decided to test his musical protocol on existing radish of the GreenHab growing in the ground.

10. Spirulina growth (GreenHab):
Everything is working as it should. We have already eaten a few grams of spirulina and made a banana and spirulina smoothie. We are now studying the impact of the type of fertilizer on the spirulina growth.

11. Aerospring (GreenHab):
Nothing to report.

12. Water consumption monitoring:
Nothing to report.

Science Report – February 21st


Crew 222 Science Report

Author : Nina SEDBON

Title : How experiments are going

– LOAC : Everything works just fine. We just took the LOAC back for the next day (or days ?) because of the rain expected. We are sending the first results tomorrow.

– MegaAres : We sent the first measures and it is coherent. That is a good start.

– Music for Plants : GreenHab officer prepared the gelatin boxes for the experiment and planted the seeds. We divided the boxes into three samples : with stimulating music, with inhibiting music, without music. Because of the temperature, we will have to put the boxes in the Science Dome.

– AeroSpring : The roots starts sprouting out of the moss, it is very exciting ! We also put the same seeds into plastic boxes in order to make a comparison with gelatin based garden and moss garden. We found out that the GreenHab was actually to hot for the gelatin to solidify. We have to put them in the Science Dome.

– The LCVG : All the fixing are almost done, espcially the sewing ones that took more time than expected.

– Water Consumption : We received the screen. The system works, we are oing to put in place in the next fex days.

– Human Factors: The entire crew takes a lot of pleasures with the experiments. We are actually playing games together. It has a positive impact on the group.

Science Report – February 17th


Crew 222 Science Report 17Feb2020
Author : Nina SEDBON
Title : Several experiments begin

In the Morning :

• LOAC and Weather Station: Put in place during the EVA. They were working just fine.
• For the LOAC the conditions were easy to obtain : stable floor, parallel to the floor for the wind captor.
• For the weather station : Girouette North orientated, three feet at a 90° angle from the ground for the whole to be very straight and vertical.

In the Afternoon:

• Aerospring : the conditions are very different in the GreenHab and in the Science Dome. We need more information to start building it because it will be very difficult to change its location once it is built. We only planted : Swiss chard, Radish, Spinach and Peas to transfer it in the vertical garden after it starts growing.
• MegaAres: Few tries in the Lower Deck to prepare the EVA for tomorrow.
• In the GreenHab : Start of the Spiruline experiment in the two aquariums.
• Backpack : Few operations in the RAM to connect the LCVG to the backpack.
• In the RAM : Fixing of the hygrometric captor for the Weather Station.
• Human Factors experiment called Foreign Languages at 16h45. It is organized with sessions of two people for 15 minutes each time. The first session is in Spanish. The two others are in English. It is a cooperation game with LEGO.
• Water Experiment: Few checks for the electrical setup. We hope to start the experiment tomorrow night. We will describe the set up once it is put in place.

Problems faced today concerning the environment of the experiments: The lack of screen for the water experiment was unexpected. A computer has to be
used to replace it.
The lack of « oscillator » for the MegaAres. It is thus complicated to test it before the EVA.

Science Report – February 14th

Crew 221 Science Report 14-FEB-20
Crew Science Officer: Hemani Kalucha

1. Microbiology Research
Today, we tested all our samples for pH and salinity.

2. ISRU project
Corn was Harvested.

3. Mung Beans

Nothing Further to report.

4. Radio Communication

Nothing further to report.

5. Geology Tests

Our geology kit was only found today in Shannon’s trailer. We conducted preliminary HCl tests on some samples.

Science Report – February 11th

Crew 221 Science Report 11-FEB-20
Crew Science Officer: Hemani Kalucha

1. Microbiology Research
Today, we completed our sample catalogue: we identified the probable bacteria species for each sample that showed growth. One of the interesting samples was the bacteria growth from a sample taken from the Candor Chasma canyon floor. The bacteria is called Flavobacterium Capsolatum.

2. ISRU project
Today, corn sprouted from the half earth soil half grey martian soil today!

3. Mung Beans

Nothing Further to report.

4. Radio Communication

Today, we disassembled the radio and Crew Engineer Louis and Crew Astronomer Ghanim reassembled the radio on an EVA, proving that it is compatibly as a mobile communication device!